A pump ejects 12 KG of water at speed of 4 metre per second in 40 seconds. find the average rate at which the pump is working

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Harsh
 
F = m*v/t
A = Fs
S = v*t
A = m*v²
Rate of working P = A/t
P = m*v²/t
P = 12000* 4²/40 = 4800 watt = 4.8 KW
average rate of working is 4.8 kW
 
Regards
Arun (askIITians forum expert)